How many ml of 4% and 1% solution are needed to make 1800 ml of 2% solution?
Answer by natolino_2017(77)

Let x: ml. of the 4% solution
y: ml. of the 1% solution
x+y = 1,800 (total solution)
0.04x +0.01y = 0,02(1,800) = 36 (total component)
Solving x=600 ml. and y= 1,200 ml.
So answer is we need 600 ml of the 4% and 1,200 ml of the 1% solution.
In a more general case, we need 1/3 of the 4% solution and 2/3 of the 1% solution to make a solution of 2%.
4%x/3+ 1%*2x/3 = 2%*x when x is the total ml. of the solution.
@natolino_